[figure: 9.7 cm]

MAGNETIC CIRCULAR POLARISATION OF SECONDARY EMISSION SPECTRA IN SINGLE CRYSTALS Y_{3}Al_{5}0_{12} : R^{3+}

(R = Tb, Dy, Er, Ho)

Valiev U.V.^{∗}, Sokolov B.Yu.^{∗}, Nekvasil V.^{∗∗} ∗Tashkent State University, 700174, Tashkent, Uzbekistan ∗∗Institute of Physics, Academy of Sciences of Czech Republic, Prague-6, 16253, Czechia

Magnetocircular polarization of luminescence (MCPL) in yttrium-aluminum garnet Y_{3}Al_{5}0_{12} containing threevalent rare-earth (RE)ions Tb, Dy, Er, Ho was investigated in the range of wavelength 450-650 nm for temperatures T =85-300K. MCPL, i.e. circular anisotropy of secondary emission spectra in a magnetic field H caused by polarization of Zeeman components of luminescence lines correspondingly to the radiative 4f - 4f transition appearing in Y_{3}Al_{5}0_{12} : R^{3+} garnets. It is discovered that, MCPL degree P = I_{+} − I_{−}/I_{+} + I_{−} , where I_{±} are the intensities of clockwise and counterclockwise polarized luminescence components, for a number radiative transitions can reach high values (P ∼ 50−60% at T =85K, H=10kOe). The behavior of MCPL degree spectra has been interpreted in view of microscopic theory which takes into account the structure of the energy spectrum for RE ion in the garnet structure, and mechanism of ”mixing” of electronic states of RE ion in external magnetic field. A possibility of using the effects fot modulation of the threshold generation of lasers to Y_{3}Al_{5}0_{12} : R^{3+}.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : Poster

Corresponding author : ?

Address for correspondence : Uzbekistan, 700011, Tashkent, Navoi st.,26/6

Email address : gleb@iaph.silk.org